find the surface area and volume of a cube whose edge is 7 cm
1
Expert's answer
2013-06-05T08:36:32-0400
Volume of a cube is equal to a^3, where a - edge of a cube. In ourcase we have a=7cm, therefore volume = 7^3=343 cm^3. The surface of a cube consists of 6 squares with side7. Then surface area = 6*( 7*7) cm^2=294 cm^2